# MINT: Multi-Vector Search Index Tuning

### v1

@@ Abstract.
Vector search plays a crucial role in many real-world applications. In addition to single-vector search, multi-vector search becomes important for multi-modal and multi-feature scenarios today. In a multi-vector database, each row is an item, each column represents a feature of items, and each cell is a high-dimensional vector. In multi-vector databases, the choice of indexes can have a significant impact on performance. Although index tuning for relational databases has been extensively studied, index tuning for multi-vector search remains unclear and challenging. In this paper, we define multi-vector search index tuning and propose a framework to solve it. Specifically, given a multi-vector search workload, we develop algorithms to find indexes that minimize latency and meet storage and recall constraints. Compared to the baseline, our latency achieves 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 × speedup.
1. Introduction
@@
Vector similarity search has become increasingly important in real-world applications, such as recommendation systems, semantic search, and retrieval-augmented generation (RAG) (Huang et al., 2020; Asai et al., 2023; Jing et al., 2024; Pan et al., 2024a). However, due to the high dimensionality and large volume of data involved, performing exact similarity search is often computationally intractable. To overcome this challenge, efficient vector search indexes have been developed to approximate search results at high accuracy (Malkov and Yashunin, 2018; Subramanya et al., 2019; Chen et al., 2021; Liu et al., 2014; Douze et al., 2024), known as Approximate Nearest Neighbor (ANN) search.
@@
Many existing approaches focus on single-domain data, overlooking that real-world datasets often contain information across multiple modalities, such as text, image, and video. For example, consider a product search on an e-commerce platform, where a user looks for “100% cotton T-shirt” with a reference picture of a white graphic T-shirt attached. If the search system relies only on text, it may miss visual details like the white color or the graphic on the T-shirt, which are essential for capturing the user’s preference for the T-shirt’s style. On the other hand, relying solely on image might neglect crucial details such as “100% cotton”, which are not apparent visually. By integrating both text and image data, a multi-modal search system can capture the user preference holistically, leading to more accurate and relevant search results. Recent research has shown that incorporating information from different modalities can significantly improve the accuracy of retrieval systems (Baltrušaitis et al., 2018; Wang et al., 2024; Duong et al., 2021; Tautkute et al., 2017).
@@
There have been many studies of multi-vector search over multi-modal information (Pan et al., 2024a; Wang et al., 2021b; Zhang et al., 2023; Chen et al., 2024; Wang et al., 2024). Each item is represented by multiple vectors, each of which is an embedding from a modality or a feature. In a database of items, each row is one item, each column is one feature, and each cell is a high-dimensional vector. A “multi-vector search” query is represented by several vectors, each of which corresponds to a feature. The query can be on all columns (i.e., features) or a subset of columns. A score function, defined between a query and an item, is an aggregation of feature-level scores, which are based on vector similarity, such as dot product or cosine distance. The query result is a set of top-$k$ items with maximal scores.
@@
One approach to answer multi-vector queries is “one-index-per-query” (Pan et al., 2024a, b). For example, let $q^{1,3}$ be a query on column-$1$ and column-$3$. This approach creates an ANN index $x^{1,3}$ on both columns. Index $x^{1,3}$ answers $q^{1,3}$ efficiently, but cannot be used to fully answer another query, say $q^{1,2,3}$ on column-$1$, column-$2$, and column-$3$. Therefore, one index must be created for each query, which consumes a lot of storage space and has long construction time. When the number of distinct column combinations on which queries are executed grows, this approach can become costly.
@@
The second approach is “one-index-per-column”. It creates several single-column indexes and uses multiple indexes to answer each query (Wang et al., 2021b; Zhang et al., 2023). In the above example of $q^{1,3}$ and $q^{1,2,3}$, one-index-per-column will create three indexes: $x^{1}$, $x^{2}$, and $x^{3}$. It answers $q^{1,3}$ by extracting items from $x^{1}$ and $x^{3}$ and then re-ranking all items together. Similarly, it answers $q^{1,2,3}$ using $x^{1}$, $x^{2}$, and $x^{3}$ together. The number of indexes is limited by the number of columns, so the storage space is constrained. However, more work is needed to evaluate queries, since the nearest neighbor of $q^{1,3}$ may not be a nearest neighbor in $x^{1}$ or $x^{3}$. Thus many candidate neighbors must be extracted from $x^{1}$ and $x^{3}$ to achieve acceptable recall for $q^{1,3}$.
@@
Using this new query execution strategy, we develop a novel algorithm for choosing the set of indexes (called a configuration) and a novel query planner. Similar to the classical index tuning in relational databases (Chaudhuri and Narasayya, 1997), given a query workload and constraints for recall and storage, the configuration searcher explores the space of possible indexes to choose a configuration that minimizes the workload’s latency while meeting the recall and storage constraints. The configuration searcher performs what-if query planning (Chaudhuri and Narasayya, 1998); for each query, it asks for the best query plan for a given set of indexes. These indexes are hypothetical indexes based on sampling, which are much less expensive than creating real indexes. Like a relational database query optimizer, the query planner uses a cost model. Given a query, it decides how best to use the hypothetical indexes to generate an efficient query execution plan.
In summary, given a set of multi-vector queries, we study how to choose which multi-column indexes to create and how to use them to answer queries. Our contributions are: (1) We define the Multi-vector Search Index Tuning problem and its building block, the Query Planning problem. We prove they are NP-hard and propose algorithms to solve them. (2) To solve these problems, we propose simplified models to estimate query latency and recall. (3) We implement Mint, a flexible index tuning framework for multi-vector search. It works with pluggable modules that estimate index performance. We demonstrate its flexibility using two types of vector index, HNSW (Malkov and Yashunin, 2018) and DiskANN (Subramanya et al., 2019). (4) We conduct experiments to measure the performance of Mint. Mint recommends indexes on millions of vectors within minutes, and its index quality outperforms the baseline by 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 ×.
This paper is organized as follows: We define the problem and provide an overview in Section 2. We present the two major modules, Query Planner and Configuration Searcher, in Section 3 and Section 4. We evaluate Mint in Section 5. Finally, we study related work in Section 6 and conclude in Section 7.
2. Problem Definition and Overview
In this section, we present the Multi-Vector Search Query and the Multi-Vector Search Index Tuning problem, challenges in solving the problem, and an overview of our solution.
We define the problem following the literature (Pan et al., 2024a, b; Wang et al., 2021b; Zhang et al., 2023).
@@ 2. Problem Definition and Overview
Consider a database of items that represents each item using three vectors. A query $q$ is on the 1st and 3rd vectors (i.e., $q.vid=\{1,3\}$). If the score function $score(\cdot,\cdot)$ is dot product, then query $q$ finds the top-$k$ items that maximize $\sum_{i\in q.vid}score(q.\textbf{v}_{i},r.\textbf{v}_{i})=q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}+q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$.
Instead of brute-force scanning of all vectors, we can create indexes to answer a query more efficiently. For instance, HNSW (Malkov and Yashunin, 2018) indexes can be created on the first ($\textbf{v}_{1}$) and the third ($\textbf{v}_{3}$) column. Then we can retrieve top results from each index based on partial score ($q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}$) and ($q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$), re-rank them using full score ($q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}+q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$), and return the top-$k$ results to the user (Wang et al., 2021b; Zhang et al., 2023). Or we can create an index on the concatenation of the first and third columns, and directly retrieve top-$k$ results. Different choices of indexes can lead to different storage footprints and different query execution plans, with different latency and recall.
@@
(1) $\min_{X}\quad$ $\sum_{(q_{i},p_{i})\in W}p_{i}\cdot latency(q_{i},X)$
(2) $s.t.\quad$ $\forall(q_{i},p_{i})\in W,recall(q_{i},X)\geq\theta_{recall}$
(3) $storage(X)\leq\theta_{storage}$
@@
We use superscripts to indicate the columns of a query or index. For example, $q^{1,3}$ is a query on the 1st and the 3rd columns, and $x^{1,3}$ is an index on the 1st and the 3rd columns. We use the term configuration to mean a set of indexes (Chaudhuri and Narasayya, 2007).
@@
Assume a workload $W=\{(q^{1,3},0.35),\allowbreak(q^{2},0.35),\allowbreak(q^{1,2,3},0.3)\}$ where $q^{1,3}$ is on the 1st and 3rd columns, $q^{2}$ is on the 2nd column, and $q^{1,2,3}$ is on all three columns. One possible configuration is $X=\{x^{1,3},x^{2}\}$, where $x^{1,3}$ may answer $q^{1,3}$, $x^{2}$ may answer $q^{2}$, and $x^{1,3}$ together with $x^{2}$ may answer $q^{1,2,3}$.
2.1. Challenges
@@
    Finding the best query plan using given indexes. When we explore the space of indexes and find a set of indexes for a given query, which subset of indexes should be used to answer the query to minimize the latency while meeting the recall threshold? We prove this problem is NP-hard in Section 3.2. Since trying all possible subsets is too expensive, we need an algorithm to quickly find a good plan for using indexes. We implement a query planner, whose role is conceptually similar to a query optimizer in relational databases but uses a completely different algorithm (Section 3.3). The query planner takes a query and a configuration as input, and outputs a query plan with the estimated latency and recall.
  2. (2)
@@
    Estimating latency and recall using simplified models. The optimization problem in Problem 1 requires good estimates of latency (Formula 1) and recall (Formula 2) for each query. Latency has a large variance across different memory, CPU, GPU, etc. Therefore, we define cost as a proxy for latency. Cost is based on the number of distance computations (Pan et al., 2024a; Malkov and Yashunin, 2018; Wei et al., 2020; Zhang et al., 2023).
@@
    Many popular indexes, such as HNSW (Malkov and Yashunin, 2018) and DiskANN (Subramanya et al., 2019), do not have a closed-form solution for their costs. Thus, index creation is probably unavoidable for estimating the cost of answering queries. Naïvely creating indexes on all items for all plans will take too long and consume too much space. For example, most graph-based indexes need hundreds-of-seconds to create for one million items (Wang et al., 2021a). Even worse, the number of indexes can be exponential in the number of columns and number of plans to consider. To solve these issues, we borrow the idea of sampling items (Chaudhuri et al., 1998) to limit the creation time of each index, and use single-column indexes to estimate the cost of multi-column indexes without creating them.
@@
    Similarly, we apply these techniques to estimate the recall, which also has no closed-form solution for many index types, including HNSW and DiskANN. More detail is in Section 3.3.2.
  3. (3)
@@
    Finding the best configuration. After we find the best query plan and estimate latency and recall for each individual query, the remaining challenge is to find the best configuration for the entire workload. Ideally, we would like to include the best indexes for all queries. However, this ideal configuration usually exceeds the storage threshold (Formula 3). Therefore, we must pick a smaller set of indexes that meets the recall and storage thresholds while minimizing cost (i.e., latency). This problem proves to be NP-hard as we will see in Section 4.1. So in Section 4.2, we develop a search algorithm that finds configurations. It outperform the baseline by 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 ×.
@@
Figure 1. Overview of Mint. The Configuration Searcher searches for the workload’s best configuration by interacting with the Query Planner and the Storage Estimator. The Query Planner finds each individual query’s best plan using the Cost Estimator and the Recall Estimator.
2.2. Overview
@@
The Storage Estimator in Figure 1 is straightforward for many types of index, including partition-based, tree-based, and graph-based. For instance, in graph-based indexes like HNSW and DiskANN, each item is a node with a fixed maximum degree, i.e., number of edges. In this case, the index’s storage usage is the number of items times the degree and the size of each edge. Next, we will focus on Query Planner and Configuration Searcher.
3. Query Planner
In this section, we define the Query Planning problem (Section 3.1), prove it is NP hard (Section 3.2), and finally develop Dynamic Programming algorithms to solve it (Section 3.3).
3.1. Definition of Query Planning
@@
Figure 2. An example query plan retrieves $k^{1}$ and $k^{3}$ items from $x^{1}$ and $x^{3}$ and then re-rank.
@@
$k^{1}$ and $k^{3}$ are parameters of the query plan. Different parameters lead to different plans with different costs. We refer to $k^{1}$ and $k^{3}$ as “extended-$k$”, denoted $ek$, for individual index scans to distinguish them from the given $k$.
A query plan can be viewed as $(X,EK)$ where $X=\{x_{i}\}$ is the indexes and $EK=\{ek_{i}\}$ is the number of items retrieved. It means that the top-$ek_{i}$ items are retrieved from index $x_{i}$ and finally re-ranked. Intuitively, a greater $ek_{i}$ means more items are retrieved, yielding a higher cost and higher recall. On the other hand, $ek_{i}=0$ means index $x_{i}$ is not used and incurs no cost.
@@
(4) $cost_{plan}(q,X,EK)=\left(\sum_{i=1}^{|X|}cost_{idx}(q,x_{i},ek_{i})\right)+cost_{rerank}(q,EK)$
@@
(5) $cost_{idx}(q,x_{i},ek_{i})=x_{i}.dim\cdot numDist(q,x_{i},ek_{i})$
@@
(6) $cost_{rerank}(q,EK)=q.dim\cdot\sum_{i=1}^{|EK|}ek_{i}$
@@
(7) $recall(q,X,EK)=\frac{\left|gt(q)\cap\left(\cup_{i=1}^{|X|}res(q,x_{i},ek_{i})\right)\right|}{|gt(q)|}$
@@
3.2. NP-Hardness of Query Planning

### v2

@@ Abstract
Vector search plays a crucial role in many real-world applications. In addition to single-vector search, multi-vector search becomes important for multi-modal and multi-feature scenarios today. In a multi-vector database, each row is an item, each column represents a feature of items, and each cell is a high-dimensional vector. In multi-vector databases, the choice of indexes can significantly impact the performance of vector search. Although index tuning for relational databases has been extensively studied, index tuning for multi-vector search remains unclear and challenging. In this paper, we define multi-vector search index tuning and propose a framework to solve it. Specifically, given a multi-vector search workload, we develop algorithms to find indexes that minimize latency and meet storage and recall constraints. Compared to the baseline, our techniques achieve a $2.1\times$ to $8.3\times$ speedup in latency.
I Introduction
@@
Vector similarity search has become increasingly important in real-world applications, such as recommendation systems, semantic search, and retrieval-augmented generation (RAG) [42, 4, 47, 63]. However, due to the high dimensionality and large volume of data involved, performing exact similarity search is often computationally intractable. To overcome this challenge, efficient vector search indexes have been developed to approximate search results at high accuracy [55, 44, 21, 52, 28], known as Approximate Nearest Neighbor (ANN) search.
@@
Many existing approaches focus on single-domain data, overlooking that real-world datasets often contain information across multiple modalities, such as text, image, and video. For example, consider a product search on an e-commerce platform, where a user looks for “100% cotton T-shirt” with a reference picture of a white graphic T-shirt attached. If the search system relies only on text, it may miss visual details like the white color or the graphic on the T-shirt, which are essential for capturing the user’s preference for the T-shirt’s style. On the other hand, relying solely on the image might overlook crucial details, such as “100% cotton”, that are not apparent from the image. By integrating both text and image data, a multi-modal search system can capture the user preference more holistically, leading to more accurate and relevant search results. Recent research has shown that incorporating information from different modalities can significantly improve the accuracy of retrieval systems [9, 78, 29, 75, 27].
@@
There have been many studies of multi-vector search over multi-modal information [63, 77, 86, 22, 78]. Each item is represented by multiple vectors, each of which is an embedding from a modality or a feature. In a database of items, each row is one item, each column is one feature, and each cell is a high-dimensional vector. A “multi-vector search” query is represented by several vectors, each of which corresponds to a feature. The query can be on all columns (i.e., features) or a subset of columns. A score function, defined between a query and an item, is an aggregation of feature-level scores, which are based on vector similarity, such as dot product or cosine distance. The query result is a set of top-$k$ items with maximal scores.
@@
One approach to answer multi-vector queries is “one-index-per-query” [63, 64]. For example, let $q^{1,3}$ be a query on column-$1$ and column-$3$. This approach creates an ANN index $x^{1,3}$ on both columns. Index $x^{1,3}$ answers $q^{1,3}$ efficiently, but cannot be used to fully answer another query, say $q^{1,2,3}$ on column-$1$, column-22, and column-$3$. Therefore, one index must be created for each query, which consumes a lot of storage space and has long construction time. When the number of distinct column combinations on which queries are executed grows, this approach can become costly.
@@
The second approach is “one-index-per-column”. It creates several single-column indexes and uses multiple indexes to answer each query [77, 86]. In the above example of $q^{1,3}$ and $q^{1,2,3}$, one-index-per-column will create three indexes: $x^{1}$, $x^{2}$, and $x^{3}$. It answers $q^{1,3}$ by extracting items from $x^{1}$ and $x^{3}$ and then re-ranking all items together. Similarly, it answers $q^{1,2,3}$ using $x^{1}$, $x^{2}$, and $x^{3}$ together. The number of indexes is limited by the number of columns, so the storage space is constrained. However, more work is needed to evaluate queries, since the nearest neighbor of $q^{1,3}$ may not be a nearest neighbor in $x^{1}$ or $x^{3}$. Thus many candidate neighbors must be extracted from $x^{1}$ and $x^{3}$ to achieve acceptable recall for $q^{1,3}$.
@@
Using this new query execution strategy, we develop a novel algorithm for choosing the set of indexes (called a configuration) and a novel query planner. Similar to the classical index tuning in relational databases [16], given a query workload and constraints for recall and storage, the configuration searcher explores the space of possible indexes to choose a configuration that minimizes the workload’s latency while meeting the recall and storage constraints. The configuration searcher performs what-if query planning [17]; for each query, it asks for the best query plan for a given set of indexes. These indexes are hypothetical indexes based on sampling, which are much less expensive than creating real indexes. Like a relational database query optimizer, the query planner uses a cost model. Given a query, it decides how best to use the hypothetical indexes to generate an efficient query execution plan.
In summary, given a set of multi-vector queries, we study how to choose which multi-column indexes to create and how to use them to answer queries. Our contributions are: (1) We define the Multi-vector Search Index Tuning problem and its building block, the Query Planning problem. We prove they are NP-hard and propose algorithms to solve them. (2) To solve these problems, we propose simplified models to estimate query latency and recall. (3) We implement Mint, a flexible index tuning framework for multi-vector search. It works with pluggable modules that estimate index performance. We demonstrate its flexibility using two types of vector index, HNSW [55] and DiskANN [44]. (4) We conduct experiments to measure the performance of Mint. Mint recommends indexes on millions of vectors within minutes, and it achieves a $2.1\times$ to $8.3\times$ speedup over the baseline.
This paper is organized as follows: We define the problem and provide an overview in Section II. We present the two major modules, Query Planner and Configuration Searcher, in Section III and Section IV. We evaluate Mint in Section V. Finally, we study related work in Section VI and conclude in Section VII.
II Problem Definition and Overview
In this section, we present the Multi-Vector Search Query and the Multi-Vector Search Index Tuning problem, challenges in solving the problem, and an overview of our solution. We define the problem following the literature [63, 64, 77, 86].
@@ II Problem Definition and Overview
Consider a database of items (e.g., e-commerce products) that represents each item using three vectors (e.g., product title, description, and main image). A query $q$ is on the 1st and 3rd vectors (i.e., $q.vid=\{1,3\}$ on product title and main image). If the score function $score(\cdot,\cdot)$ is dot product, then query $q$ finds the top-$k$ items that maximize $\sum_{i\in q.vid}score(q.\textbf{v}_{i},r.\textbf{v}_{i})=q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}+q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$.
Instead of brute-force scanning of all vectors, we can create indexes to answer a query more efficiently. For instance, HNSW [55] indexes can be created on the first ($\textbf{v}_{1}$) and the third ($\textbf{v}_{3}$) column. Then we can retrieve top results from each index based on partial score ($q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}$) and ($q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$), re-rank them using full score ($q.\textbf{v}_{1}\cdot r.\textbf{v}_{1}+q.\textbf{v}_{3}\cdot r.\textbf{v}_{3}$), and return the top-$k$ results to the user [77, 86]. Or we can create an index on the concatenation of the first and third columns, and directly retrieve top-$k$ results. Different choices of indexes can lead to different storage footprints and different query execution plans, with different latency and recall.
@@
$\min_{X}\quad$ $\sum_{(q_{i},p_{i})\in W}p_{i}\cdot latency(q_{i},X)$ (1)
$s.t.\quad$ $\forall(q_{i},p_{i})\in W,recall(q_{i},X)\geq\theta_{recall}$ (2)
$storage(X)\leq\theta_{storage}$ (3)
@@
We use superscripts to indicate the columns of a query or index. For example, $q^{1,3}$ is a query on the 1st and the 3rd columns, and $x^{1,3}$ is an index on the 1st and the 3rd columns. We use the term configuration to mean a set of indexes [18].
@@
Assume a workload $W=\{(q^{1,3},0.35),\allowbreak(q^{2},0.35),\allowbreak(q^{1,2,3},0.3)\}$ where $q^{1,3}$ is on the 1st and 3rd columns (e.g., product title and main image), $q^{2}$ is on the 2nd column (e.g., description), and $q^{1,2,3}$ is on all three columns. If we are allowed to create only two indexes, a possible configuration is $X=\{x^{1,3},x^{2}\}$, where $x^{1,3}$ may answer $q^{1,3}$, $x^{2}$ may answer $q^{2}$, and $x^{1,3}$ together with $x^{2}$ may answer $q^{1,2,3}$.
II-A Challenges
@@
    Finding the best query plan using given indexes. When we explore the space of indexes and find a set of indexes for a given query, which subset of indexes should be used to answer the query to minimize the latency while meeting the recall threshold? We prove this problem is NP-hard in Section III-B. Since trying all possible subsets is too expensive, we need an algorithm to quickly find a good plan for using indexes. We implement a query planner, whose role is conceptually similar to a query optimizer in relational databases but uses a completely different algorithm (Section III-C). The query planner takes a query and a configuration as input, and outputs a query plan with the estimated latency and recall.
  2. 2.
@@
    Estimating latency and recall using simplified models. The optimization problem in Problem 1 requires good estimates of latency (Formula 1) and recall (Formula 2) for each query. Latency has a large variance across different memory, CPU, GPU, etc. Therefore, we define cost as a proxy for latency. Cost is based on the number of distance computations [63, 55, 80, 86].
@@
    Many popular indexes, such as HNSW [55] and DiskANN [44], do not have a closed-form solution for their costs. Thus, index creation is probably unavoidable for estimating the cost of answering queries. Naïvely creating indexes on all items for all plans will take too long and consume too much space. For example, most graph-based indexes need hundreds of seconds to create for one million items [79]. Even worse, the number of indexes can be exponential in the number of columns and number of plans to consider. To solve these issues, we borrow the idea of sampling items [15] to limit the creation time of each index, and use single-column indexes to estimate the cost of multi-column indexes without creating them.
@@
    Similarly, we apply these techniques to estimate the recall, which also has no closed-form solution for many index types, including HNSW and DiskANN. More detail is in Section III-C2.
    In theory, one can always construct an adversarial query requiring a linear scan of $O(n)$ vectors (i.e., expensive cost) to achieve an acceptable recall, where $n$ is the total number of rows [43], but in practice, our estimations work reasonably well as we show in experiments (Section V).
  3. 3.
@@
    Finding the best configuration. After we find the best query plan and estimate latency and recall for each individual query, the remaining challenge is to find the best configuration for the entire workload. Ideally, we would like to include the best indexes for all queries. However, this ideal configuration usually exceeds the storage threshold (Formula 3). Therefore, we must pick a smaller set of indexes that meets the recall and storage thresholds while minimizing cost (i.e., latency). This problem proves to be NP-hard as we will see in Section IV-A. So in Section IV-B, we develop a search algorithm that finds configurations. It outperforms the baseline by $2.1\times$ to $8.3\times$.
@@
Figure 1: Overview of Mint. The Configuration Searcher searches for the workload’s best configuration by interacting with the Query Planner and the Storage Estimator. The Query Planner finds each individual query’s best plan using the Cost Estimator and the Recall Estimator.
II-B Overview
@@
The Storage Estimator in Figure 1 is straightforward for many types of index, including partition-based, tree-based, and graph-based. For instance, in graph-based indexes like HNSW and DiskANN, each item is a node with a fixed maximum degree, i.e., number of edges. In this case, the index’s storage usage is the number of items times the degree and the size of each edge. Note that each index only stores the graph using item IDs without duplicating the underlying vectors. So the index graph size is independent of the dimensions of vectors. [^1] For some index techniques like LSH (which needs more than linear space [57, 2]) and subspace collision framework [81], the index storage is not necessarily linear in the number of items. In such cases, technique-specific storage estimators are reqruied and can be plugged into our framework. Next, we will focus on Query Planner and Configuration Searcher.
III Query Planner
Next, we define the Query Planning problem (Section III-A), prove it is NP-hard (Section III-B), and develop Dynamic Programming algorithms to solve it (Section III-C).
III-A Definition of Query Planning
@@
Figure 2: An example query plan retrieves $k^{1}$ and $k^{3}$ items from $x^{1}$ and $x^{3}$ and then re-rank.
@@
$k^{1}$ and $k^{3}$ are parameters of the query plan. Different parameters lead to different plans with different costs. We refer to $k^{1}$ and $k^{3}$ as “extended-kk”, denoted $ek$, for individual index scans to distinguish them from the given $k$.
A query plan can be viewed as $(X,EK)$ where $X=\{x_{i}\}$ is the indexes and $EK=\{ek_{i}\}$ is the number of items retrieved. It means that the top-$ek_{i}$ items are retrieved from index $x_{i}$ and finally re-ranked. Intuitively, a greater $ek_{i}$ means more items are retrieved, yielding a higher cost and higher recall, while $ek_{i}=0$ means index $x_{i}$ is not used and incurs no cost.
@@
$cost_{plan}(q,X,EK)=\left(\sum_{i=1}^{|X|}cost_{idx}(q,x_{i},ek_{i})\right)+cost_{rerank}(q,EK)$ (4)
@@
$cost_{idx}(q,x_{i},ek_{i})=x_{i}.dim\cdot numDist(q,x_{i},ek_{i})$ (5)
@@
$cost_{rerank}(q,EK)=q.dim\cdot\sum_{i=1}^{|EK|}ek_{i}$ (6)
@@
$recall(q,X,EK)=\frac{\left|gt(q)\cap\left(\cup_{i=1}^{|X|}res(q,x_{i},ek_{i})\right)\right|}{|gt(q)|}$ (7)
@@
III-B NP-Hardness of Query Planning